OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

I would suggest finding the slope using the slope formula first.

OpenStudy (anonymous):

recall y=mx+b so solve the following system 2=-2m+b 1=5m+b

OpenStudy (anonymous):

that was helping helper michael

OpenStudy (anonymous):

Im sorry Im not understanding this problem

OpenStudy (anonymous):

I would actually use \[m=\frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

Use that ^^^ to find the slope first.

OpenStudy (anonymous):

m = 1-2 5-2

OpenStudy (anonymous):

-1-3

OpenStudy (anonymous):

Everything is correct except not 5 - 2 (remember that E is (-2,2))

OpenStudy (anonymous):

oh ok so -1- -7

OpenStudy (anonymous):

Im still lost Paynesdad

OpenStudy (anonymous):

Ok so ...\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }= \frac{ 1-2 }{ 5-(-2)} \] ... you had 5- 2 instead of 5 - (-2)... don't lose those negatives!

OpenStudy (anonymous):

y=-1 x=-7

OpenStudy (anonymous):

5 - (-2) is not -7

OpenStudy (anonymous):

x=7 sorry

OpenStudy (anonymous):

So you have m = -1/7.

OpenStudy (anonymous):

Now what is the point slope formula?

OpenStudy (anonymous):

so the answer is 7x- (-1) y

OpenStudy (anonymous):

what is the point slope formula

OpenStudy (anonymous):

\[y - y _{1}=m \left( x-x _{1} \right)\]